Suppose that a firm’s cost function is given by: C(q)=4q. If the market price is $10, then the firm’s profit from producing and selling 200 units is?

$1100

$1000

$1300

$1200

$1400
Answer #1

Answer;

Profit = P*Q – C(q)

Given: P = $10

Q = 200 units

C(q) = 4q

Hence profit = 10 * 200 – 4(200) = 2000 – 800 = 1200

Profit = $1200

Correct option: D] $1200